looking for some help and need it too. one question has me stoped dead and two others just want someone to check my work thank you.
Among 100 people randomly chosen in a survey, 10 stated that they believed that the elected representatives are honest most of the time. The approximate 95% confidence interval for the percentage of the population who believe that such people are honest most of the time is
For questions 11 and 12 use the table immediately below. The table shows the number of smokers and nonsmokers by gender among 210 randomly selected people.

Smokers Nonsmokers Totals
Male 30 60 90
Female 40 80 120
Totals 70 140 210

_____11. If a smoker is chosen at random, the probability that the person is male is

12. If a male is chosen at random, the probability that he is a smoker is
would like to get a nother look at 11 and 12 I think 11 is 30/70 and 12 is no answer with this data.
thank you for your time.

Answer by stanbon(75887)

Among 100 people randomly chosen in a survey, 10 stated that they believed that the elected representatives are honest most of the time. The approximate 95% confidence interval for the percentage of the population who believe that such people are honest most of the time is
------
E=z*sqrt[(pq)/n]
E=1.96sqrt[[(1/10*9/10)/100]
E=1.96sqrt[9/10000]
E=1.96*3/100
E=0.0588
----------
p-hat=10/100= 0.1
CI: 0.1-0.0588 or, in interval notation, (0.412,1.0588)
=================

For questions 11 and 12 use the table immediately below. The table shows the number of smokers and nonsmokers by gender among 210 randomly selected people.
Smokers Nonsmokers Totals
Male 30 60 90
Female 40 80 120
Totals 70 140 210
_____11. If a smoker is chosen at random, the probability that the person is male is
P(male|smoker) = 30/70 = 3/7
12. If a male is chosen at random, the probability that he is a smoker is
P(smoker|male} = 30/90 = 1/3
===============
Cheers,
Stan H.